### v1

# From Spin Waves to Monte Carlo Simulations: Compiling an Experimental Exchange Interaction Dataset for Magnetic Materials

Abstract
Inelastic neutron scattering data on magnetic crystals are highly valuable in materials science, as they provide direct insight into microscopic magnetic interactions. Using spin wave theory, these interactions can be extracted from magnetic excitations observed in such experiments. However, these datasets are often scattered across the literature and lack standardization, limiting their accessibility and usability. In this paper, we compile and standardize Heisenberg exchange interaction data for magnetic materials obtained from inelastic neutron scattering experiments. Through an extensive literature review, we identify experimental data for approximately 100 magnetic materials. The standardized dataset includes mapping the results of various Heisenberg Hamiltonians into a unified standard form, visualizations of crystal structures with annotated exchange interactions, and input and output files from Monte Carlo simulations performed for each compound using the ESpinS code. Using experimentally determined exchange interactions, we calculate transition temperatures ($T_{c}$) via classical Monte Carlo simulations. Additionally, we assess the effectiveness of the $(S+1)/S$ correction within classical Monte Carlo simulations, finding that it produces transition temperatures in excellent agreement with experimental values in most cases. The complete dataset, along with supporting resources, is publicly available on GitHub.
I Introduction
Inelastic neutron scattering (INS) is a powerful technique for probing magnetic excitations, such as magnons—quasiparticles that represent collective spin-wave excitations in a lattice [1, 2, 3]. By measuring the energy and momentum transferred during neutron scattering events, INS provides direct insight into magnon dispersion relations. These dispersions are highly sensitive to the underlying magnetic exchange interactions, making INS an invaluable tool for characterizing magnetic systems.
Magnetic interactions in a material can be extracted by analyzing magnon excitation data through spin-wave theory, which models magnetic excitations based on an effective spin Hamiltonian [1, 4]. However, employing full quantum spin operators in such Hamiltonians introduces significant theoretical and computational complexity. To overcome this challenge, linear spin-wave theory (LSWT) serves as a widely used approximation method [5, 6, 1]. LSWT simplifies the problem by linearizing quantum fluctuations around an ordered magnetic ground state, expressing spin operators in terms of bosonic creation and annihilation operators. This reduces the Hamiltonian to a quadratic form, enabling an analytical solution. For example, in a simple antiferromagnetic system with only nearest-neighbor interactions, the spin Hamiltonian is given by: $\mathcal{H}=-\sum_{\langle i,j\rangle}J\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ where $J$ is the exchange interaction, and the summation $\langle i,j\rangle$ runs over nearest-neighbor spin pairs. Within LSWT, the magnon dispersion relation for this model simplifies to: $E(\mathbf{k})=-JZS\sqrt{1-\gamma_{\mathbf{k}}^{2}}$ where $Z$ is the coordination number (i.e., the number of nearest neighbors), and $\gamma_{\mathbf{k}}$ is a geometric structure factor defined as: $\gamma_{\mathbf{k}}=\frac{1}{Z}\sum_{\bm{\delta}}e^{i\mathbf{k}\cdot\bm{\delta}}$. Here, $\bm{\delta}$ represents the displacement vectors connecting a reference site to its nearest neighbors [6].
Despite the availability of theoretical tools such as LSWT for analyzing INS data, INS measurements remain limited due to the specialized and resource-intensive equipment required. Our literature review identifies only about 100 studies that report both magnon INS data and corresponding spin model mappings using LSWT. Compiling and standardizing this data into a unified framework would greatly benefit magnetism research.
Such a dataset could serve as a foundation for integrating future high-quality INS results and support systematic theoretical studies aimed at improving computational methods and predictive models. Additionally, establishing validation criteria is essential, as discrepancies often arise between studies on the same material, or even within a single study where multiple spin models are proposed. A fast, reliable theoretical approach such as classical Monte Carlo (MC) [7] to evaluate model accuracy would greatly assist in identifying the most appropriate spin Hamiltonian
When spin-wave experimental values for Heisenberg exchange interactions are directly applied in classical MC simulations, the resulting transition temperatures $T_{c}$ often deviate from expectations. These discrepancies stem from the quantum effects accounted for in spin-wave theory, which are integral to deriving Heisenberg exchange parameters from neutron scattering experiments. To reconcile this mismatch in classical MC simulations, the $(S+1)/S$ correction, where $S$ is the spin magnitude of magnetic atoms (spin quantum number), has been proposed—either directly to the exchange parameters or indirectly to the $T_{c}$ values obtained from the simulations [8, 9, 10]. In a previous study [8], we used Heisenberg exchange interactions derived from INS data for 13 magnetic materials to predict their transition temperatures using classical MC simulations. By applying the $(S+1)/S$ correction to the MC results, we achieved a mean absolute percentage error (MAPE) of 8% in the predicted transition temperatures.
Building on our previous findings, this work aims to systematically assess the influence of the $(S+1)/S$ correction on magnetic transition temperatures across a broader range of materials. We extend our dataset to encompass approximately 73 inelastic neutron scattering (INS) studies, ensuring that all extracted exchange parameters are standardized within a consistent spin model Hamiltonian. Using classical MC simulations, we compute transition temperatures both with and without the correction. Our analysis confirms that incorporating the $(S+1)/S$ factor significantly improves the agreement between simulated and experimental $T_{c}$ values. To support further research and reproducibility, the complete dataset has been made publicly available on GitHub [11].
The paper is organized as follows: the Methods section outlines the methodology employed in this study, while the Results section examines the use of classical MC simulations both for predicting magnetic transition temperatures and for identifying inaccuracies in reported exchange parameters derived from INS data. Finally, the Conclusions section summarizes our findings and discusses their broader implications.
II Method
II.1 Data gathering and standardizing
We attempted to identify as many research papers as possible that present INS data (magnetic excitations) analyzed using spin wave theory to extract exchange interactions. A review of the literature shows that the Heisenberg term in spin model Hamiltonians typically appears in one of the following forms:
$H_{Heis.}$ $=$ $-\sum_{i,j}J_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$
$H_{Heis}$ $=$ $-\frac{1}{2}\sum_{i,j}J_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$
$H_{Heis}$ $=$ $-\sum_{<i,j>}J_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$
$H_{Heis}$ $=$ $-2\sum_{<i,j>}J_{i,j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}$
Figure 1: Illustrative images of sample crystal structures from our database. Each image highlights the exchange interactions and their labeling, as presented in the corresponding INS paper. The labels in parentheses indicate the ranking of exchange interactions based on nearest-neighbor distances (e.g., $J_{1}$ corresponds to the first nearest neighbor,$J_{2}$ to the second nearest neighbor, and so on).
One of the key challenges in standardizing exchange interactions is the ambiguity in the choice of spin Hamiltonian conventions, particularly regarding the double counting of pairwise interactions. For example, models that use $-\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $-\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ are designed so that each interaction between spins at sites $i$ and $j$ is counted only once. In contrast, models that adopt $-\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $-2\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ count each pairwise interaction twice. A major source of confusion arises when some studies claim to use the convention $-\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, yet a detailed analysis—often involving comparisons with other works—reveals that their reported exchange constants have already been corrected to avoid double counting. In practice, these studies effectively use the convention $-\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, as seen, for example, in Refs. [12, 13, 14, 15, 3].
A particularly clear example of this mismatch between the stated Hamiltonian and the magnon dispersion derived from spin wave theory is found in the altermagnet MnF2 [15, 3]. This compound has been extensively studied in INS experiments [16, 17, 15, 18] and is widely regarded as a textbook prototype of antiferromagnetic order [3, 19, 20].
Additionally, some papers use a positive sign in front of the summation instead of a negative one. In such cases, the interpretation of $J_{i,j}$ is reversed: a negative value of $J_{i,j}$ indicates a ferromagnetic interaction, while a positive value corresponds to an antiferromagnetic interaction.
For magnetic anisotropy, the following forms are commonly considered in the literature:
$H_{\text{Ani}}$ $=$ $-D\sum_{i}S_{i,z}^{2}$
$H_{\text{Ani}}$ $=$ $-D_{x}\sum_{i}S_{i,x}^{2}-D_{z}\sum_{i}S_{i,z}^{2},$
where $D$, $D_{x}$, and $D_{z}$ represent the anisotropy strengths along specific directions.
To ensure consistency, we express the exchange and magnetic anisotropy interactions in the following standardized form:
$H=-\frac{1}{2}\sum_{i,j}\tilde{J}_{ij}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}+\sum_{i}\hat{{\mathbf{S}}}_{i}\tilde{{\mathbf{A}}}\hat{{\mathbf{S}}}_{i}.$ (1)
Here, $\tilde{J}_{ij}$ (obtained by rescaling ${J}_{ij}$ by $S^{2}$) represents the Heisenberg exchange interaction strength between sites $i$ and $j$, while $\hat{\mathbf{S}}_{i}$ and $\hat{\mathbf{S}}_{j}$ denote unit vectors indicating the magnetic moment direction at lattice sites $i$ and $j$. The matrix $\tilde{{\mathbf{A}}}$ characterizes the anisotropy.
To avoid additional complexity, we do not consider the Dzyaloshinskii-Moriya interaction (DMI) in this work. Therefore, we select data where DMI is either not reported or is negligible compared to the dominant Heisenberg exchange interaction.
Standardization also requires knowledge of the spin quantum number $S$, as it is necessary for rescaling $J_{ij}$ to $\tilde{J}_{ij}$. Consequently, we exclude studies where the value of $S$ is ambiguous. When available, we adopt the value of $S$ reported in INS studies, as it is typically consistent with the experimentally measured magnetic moment.
In the literature, interaction strengths are expressed in various units, including THz, meV, Kelvin, and $\text{cm}^{-1}$. More recent studies tend to report magnetic interactions in meV. Therefore, we standardize all interaction strengths in meV for consistency.
II.2 Visualization and MC simulations
For each compound, provide an image of the crystal structure including illustration for $J_{i,j}$ between first, second ,.., nth nearest neighbors of magnetic atoms indicating by $J_{1}$, $J_{2}$ and $J_{n}$ respectively (Fig. 1). In some papers, researchers prefer to indicate Heisenberg interactions $J_{i,j}$ by indices of lattice vectors (e.g, $J_{c}$ indicates exchange interaction between magnetic atoms along lattice c axis). In such cases, we indicate our standard naming of exchange interactions inside parentheses (Fig. 1).
We use VESTA [21] for structure visualization. For each compound, we provide structure files in VESTA format, which not only allow researchers to visualize the atomic positions and lattice vectors but also highlight Heisenberg exchange interactions using thick red lines.
We use the ESpinS [22] code for MC simulations. For each compound, we provide all the necessary input and output files for ESpinS code [11]. Our simulations indicate that using approximately 2000 lattice sites is generally sufficient to predict the transition temperature. Therefore, for MC simulations, we use supercells that contain a minimum of 2000 sites. To determine the transition temperature, we identify the peak of the magnetic specific heat, given by $C_{M}=\frac{\braket{E^{2}}-\braket{E}^{2}}{Nk_{B}T^{2}}$. In cases where the exact peak position is unclear, we also analyze the fourth-order energy cumulant, defined as $U_{E}=1-\frac{1}{3}\frac{\braket{E^{4}}}{\braket{E^{2}}^{2}}$.
II.3 Data on Github
For each material in the GitHub database [11], there is a dedicated directory named after the compound. Each directory contains the following information:
  • Structure file: A VESTA format file with thick lines indicating exchange interactions.
  • Visualization: A JPEG image showing atomic positions and exchange interactions, as depicted in Fig. 1. This image is displayed online for each compound.
  • Data tables: Tables listing exchange interactions (in meV), spin quantum number $S$, and both experimental and MC transition temperatures.
  • Reference links: Sources for INS data and transition temperature values.
  • MC simulation files: A MC directory containing all input and output files from MC simulations performed using ESpinS.
III Results and discussion
III.1 Predicting transition temperature using Monte Carlo
The quantum Monte Carlo (QMC) method is among the most accurate approaches for determining the transition temperatures of magnetic materials. However, its applicability is often limited by the fermion sign problem, especially in frustrated systems  [23]. Additionally, accounting for finite-size effects in QMC simulations is highly computationally demanding.
Other methods, such as mean field theory (MFT) [24] and the random phase approximation (RPA) [25, 26, 27], are also used. MFT tends to overestimate the transition temperature compared to experimental results, while RPA improves upon MFT and yields more accurate predictions [26, 27, 28, 29]. However, applying RPA requires specific techniques and details, such as the magnetic ordering vector, which limits its use to researchers with sufficient theoretical knowledge [28].
Classical MC can be a useful tool for providing a rough estimate of the transition temperature in magnetic materials. However, when using INS data fitted with spin-wave theory, classical MC simulations tend to underestimate the transition temperature [8]. This discrepancy casts doubt on the reliability of the INS data [30]. Since spin-wave theory incorporates quantum mechanical effects in fitting inelastic neutron scattering (INS) data, including the factor $(S+1)/S$ is essential for obtaining accurate results from classical Monte Carlo (MC) simulations. This factor arises from comparing the quantum and classical expressions for the expectation value of the squared spin operator. In the quantum case, $\langle\mathbf{S}^{2}\rangle$ is proportional to $S(S+1)$, whereas in the classical limit it scales as $S^{2}$. To reconcile these two approaches, either the exchange parameters used in classical MC simulations or the transition temperatures obtained from classical MC must be rescaled by the factor $(S+1)/S$ [8].
Figure 2 shows the transition temperatures obtained from MC simulations ($T_{\mathrm{MC}}$) for all 73 compounds, revealing a significant deviation from the experimental transition temperatures. However, applying the $(S+1)/S$ correction:
$T_{\mathrm{MC}}^{*}=\frac{S+1}{S}T_{\mathrm{MC}}$ (2)
significantly improves the accuracy of transition temperature predictions.
The MAPE for $T_{\mathrm{MC}}^{*}$ is 9.6%, whereas for $T_{\mathrm{MC}}$, the error is much greater, equal to 35.1%. For 37% of the compounds, the absolute percentage error (APE) is less than 5%. In 30% of the cases, APE falls between 5% and 10%, while for 12% of the compounds, it ranges from 10% to 15%. Only 20% of the compounds exhibit an APE greater than 15%. The detailed data from the MC simulations shown in Fig. 2 are provided in Appendix A.
It is important to note that exchange interactions derived from INS data are subject to uncertainties due to their reliance on pure fitting to LSWT. Additionally, exchange interactions can vary with temperature in practice [31, 32, 33]. These factors collectively impact the accuracy of predicting the transition temperature using simulation methods such as MC.
As a result, calculating an exact transition temperature is not solely dependent on theoretical methods; it also relies on the accuracy of experimentally derived exchange interactions. This means that even with a perfect theoretical approach, expecting a transition temperature that exactly matches experimental results is unrealistic.
III.2 Refining Spin Hamiltonians Using Monte Carlo Results
From the plot (Fig. 2), we can establish two key observations:
  1. 1.
    The transition temperature obtained from classical MC simulations ($T_{\mathrm{MC}}$) is always lower than the experimental transition temperature ($T_{\mathrm{Exp.}}$), i.e., $T_{\mathrm{MC}}<T_{\mathrm{Exp.}}$.
  2. 2.
    The quantum correction factor $(S+1)/S$ accurately accounts for this discrepancy in most cases.
These observations provide useful criteria for assessing the accuracy of exchange parameters derived from INS data using spin-wave theory. In the following, we discuss several cases to illustrate how these criteria can be used to identify better models and filter out unreliable data.
Figure 2: Results from MC simulations for 73 compounds. $T_{\mathrm{MC}}$ represents the transition temperature obtained from MC simulations using exchange parameters derived from INS data fitted with spin-wave theory. $T_{\mathrm{MC}}^{*}$ denotes the quantum-corrected transition temperature, obtained by applying the correction factor $(S+1)/S$ to $T_{\mathrm{MC}}$. All data points shown in this diagram are provided in Appendix A.
NiO is one of the most extensively studied antiferromagnetic compounds, with a transition temperature of 523 K and a quantum spin number $S=1$. However, only a single INS dataset exists for it [34]. In the INS study, two different Heisenberg models were used to fit the data: a $J_{1}-J_{2}$ model (including only first- and second-nearest-neighbor exchange interactions) and a $J_{1}-J_{4}$ model (including interactions up to the fourth nearest neighbor).
Using the $J_{1}-J_{2}$ model, the MC simulation yields $T_{\mathrm{MC}}=310$ K and $T_{\mathrm{MC}}^{*}=620$ K. In contrast, the $J_{1}$-$J_{4}$ model produces $T_{\mathrm{MC}}=304$ K and $T_{\mathrm{MC}}^{*}=608$ K, suggesting that the $J_{1}-J_{4}$ model provides a more accurate estimate of the transition temperature.
[…]
Another example is MnBi2Te4, for which two successive research papers were published by overlapping groups of authors [14, 35]. This compound has an experimental magnetic transition temperature of 24 K and a spin quantum number of $S=5/2$. All spin models considered in these studies include intra-layer exchange interactions ($J_{1},J_{2},\dots$) and an inter-layer exchange interaction ($J_{c}$).
[…]
During our investigation, we found instances where the transition temperature obtained from Monte Carlo simulations ($T_{\mathrm{MC}}$) exceeded the experimental value, and other cases where the quantum-corrected temperature ($T_{\mathrm{MC}}^{*}$) was significantly lower. Such discrepancies often indicate problems with the INS data or the fitting procedure. These issues may stem from an incorrectly defined Hamiltonian—such as a missing factor of $1/2$—poor data fitting, or an inaccurate estimate of the spin quantum number $S$. For example, using the INS-derived exchange interactions for LiNiPO4 [36], LiCoPO4 [37], and LiMnPO4 [38], the MC-calculated transition temperatures ($T_{\mathrm{MC}}$) are overestimated. This points to a mismatch between the spin Hamiltonian and the spin-wave formalism, likely caused by an omitted factor of $1/2$ in the Hamiltonian. A closer examination of the spin-wave equations used in these studies supports this conclusion.
IV Conclusions
Through analysis of more than 100 INS studies that derived exchange interactions from spin-wave theory, we identified 73 compounds with reliable data and compiled them into a standardized database. This database includes exchange interaction parameters, spin quantum numbers, reference sources, MC simulation results, structure files, and visualized structures. Using this dataset, we investigated the accuracy of transition temperature predictions by applying the $(S+1)/S$ correction to classical MC results. Our findings show that, in most cases, the corrected results align more closely with experimental values, achieving a mean absolute percentage error (MAPE) of 9%. We propose classical MC simulations and the $(S+1)/S$ correction as useful tools for refining INS data and identifying potential inconsistencies.
Acknowledgements.
M. A. acknowledges financial support from the Iran National Science Foundation (INSF) under Project No. 4002648. N. R. and A. R. O. acknowledge support from the Russian Science Foundation under Grant No. 19-72-30043.
Appendix A Transition temperature
In this appendix, we provide all the data required to plot Figure 2. Table A1 summarizes the experimental transition temperatures ($T_{\mathrm{Exp.}}$), Monte Carlo (MC) transition temperatures ($T_{\mathrm{MC}}$), and quantum-corrected MC transition temperatures ($T_{\mathrm{MC}^{*}}$). For each compound, we include references to the inelastic neutron scattering (INS) data along with the corresponding experimental transition temperatures. Additional details, such as the input and output files from the MC simulations, are available in the GitHub repository [11]
Table A1: Transition temperatures obtained from MC simulations. $T_{\mathrm{MC}}$ was determined through MC simulations. $T_{\mathrm{MC}}^{*}$ was then obtained by multiplying the MC results with the ($\frac{S+1}{S}$) factor. Additionally, $T_{\mathrm{Exp.}}$ represents the experimentally reported temperature. $S$ represents the spin of the magnetic ion, and Error denotes the percentage difference between $T_{\mathrm{MC}}^{*}$ and $T_{\mathrm{Exp.}}$.
$T_{\mathrm{Exp.}}$(K) $T_{\mathrm{MC}}$(K) $T_{\mathrm{MC}}^{*}$(K) S Error(%)
Ba2NiWO6[39] $48$ [39] $23$ $46$ $1$ $4.2$
BaMn2As2[40] $625$[40] $460$ $644$ $5/2$ $3.0$
BaMn2Bi2[41] $387.2$[41] $238$ $333.2$ $5/2$ $13.9$
BaNi2As2O8[42] $18.5$[43] $11$ $22$ $1$ $18.9$
Bi2CuO4[44] $50$ [44] $17.5$ $52.5$ $1/2$ $5$
BiFeO3[45] $650$[45] $431$ $603.4$ $5/2$ $7.2$
Ca3Ru2O7[46] $56$[46] $30$ $60$ $1$ $7.1$
CaMn2Sb2[47] $85$[47] $69$ $115$ $3/2$ $35.3$
CaMn7O12[48] $90$[48] $61$ $122$ $1$ $35.5$
CaMnBi2[49] $270$[49] $209$ $292.6$ $5/2$ $8.4$
CoO[50] $289$[50] $201$ $353$/2323/23 / 2 $15.9$
CoPS3[51] $120$[51] $61$ $101.7$ $3/2$ $15.3$
Cr2O3[52] $308$[52] $194$ $23.33$/2323/23 / 2 $5.0$
Cr2TeO6[53] $93$[53] $53$ $88.3$ $3/2$ $5.0$
Cr2WO6[54] $45$[54] $27$ $45$ $3/2$ $0.0$
CrBr3[55] $32$[55] $20$ $3.33$/2323/23 / 2 $4.2$
CrCl2[56] $20$[56] $15$ $22.5$ $2$ $12.5$
CrCl3[57] $14$[57] $10$ $16.6$ $3/2$ $18.6$
CrI3[58] $61$[58] $39$ $65$ $3/2$ $6.6$
CuFeO2[59] $11$[59] $7.5$ $10.5$ $5/2$ $4.5$
CuO[60] $213$[60] $72$ $216$ $1/2$ $1.4$
EuO[61] $69.15$[61] $52$ $66.9$ $7/2$ $3.3$
EuS[62] $16.57$[62] $13$ $16.7$ $7/2$ $0.8$
Fe2O3[63] $960$[63] $730$ $1022$ $5/2$ $6.4$
FeCl2[64] $23.55$[64] $11.5$ $23$ $1$ $2.3$
FeF2[65] $78.4$[65] $55$ $82.5$ $2$ $5.2$
FeO[66] $198$[66] $148$ $222$ $2$ $12.1$
FePS3[67] $120$[67] $91$ $136.5$ $2$ $13.8$
FePSe3[68] $110$[68] $75$ $112.5$ $2$ $2.3$
FePt3[69] $170$[69] $104$ $167.0$ $1.65$ $1.8$
HoFeO3[70] $647$[70] $456$ $638.4$ $5/2$ $1.3$
KCoF3[71] $114$[72] $33$ $99$ $1/2$ $13.2$
KCuF3[73] $39$[73] $15$ $45$ $1/2$ $15.4$
La2CoO4[74] $275$[74] $161$ $268.3$ $3/2$ $2.4$
La2NiO4[75] $328$[75] $220$ $440$ $1$ $34.1$
LaFeO3[76] $738$[77] $504$ $705.6$ $5/2$ $4.4$
$T_{\mathrm{Exp.}}$(K) $T_{\mathrm{MC}}$(K) $T_{\mathrm{MC}}^{*}$(K) S Error(%)
LaMnO3[78] $139.5$[78] $101$ $151.5$ $2$ $8.6$
LaVO3[76] $150$[79] $70$ $140$ $1$ $6.7$
LiFePO4[80] $50$[80] $41$ $61.5$ $2$ $23$
LuMnO3[81] $87.5$[81] $71$ $106.5$ $2$ $21.7$
MnBi[82] $630$[82] $394$ $591$ $2$ $6.2$
MnBi2Te4[35] $24$[35] $16$ $22.4$ $5/2$ $6.7$
MnF2[16] $67$[16] $48$ $67.2$ $5/2$ $0.3$
MnO[83] $117$[83] $79$ $110.6$ $5/2$ $5.5$
MnO2[84] $92$[84] $63$ $105$ $3/2$ $14.1$
MnPS3[85] $78$[85] $61$ $85.4$ $5/2$ $9.5$
MnPSe3[86] $74$[86] $52$ $72.8$ $5/2$ $1.6$
MnS[87] $152$[87] $105$ $147$ $5/2$ $3.3$
MnS2[88] $48$[88] $25$ $35$ $5/2$ $27.1$
MnTe[89] $310$[89] $251$ $351.4$ $5/2$ $13.4$
MnTiO3[90] $65$[90] $45$ $63$ $5/2$ $3.1$
MnWO4[91] $13.5$[91] $9$ $12.6$ $5/2$ $6.7$
NdFeO3[92] $693$[92] $463$ $648.2$ $5/2$ $6.5$
NiBr2[93] $52$[93] $28$ $56$ $1$ $7.7$
NiCl2[94] $52.3$[94] $36$ $72$ $1$ $37.7$
NiF2[95] $73.2$[95] $42$ $84$ $1$ $14.8$
NiO[34] $523$[34] $310$ $620$ $1$ $18.5$
PrFeO3[96] $707$[96] $471$ $659.4$ $5/2$ $6.7$
Rb2MnF4[97] $38.4$[97] $33$ $46.2$ $5/2$ $20.3$
RbMnCl3[98] $94$[98] $69$ $96.6$ $5/2$ $2.8$
RbMnF3[99] $82.6$[99] $54$ $75.6$ $5/2$ $8.5$
RbNiF3[100] $133$[100] $79$ $158$ $1$ $18.8$
Sr2NiWO6[101] $54$[101] $26$ $52$ $1$ $3.7$
SrMnBi2[102] $290$[102] $219$ $306.6$ $5/2$ $5.7$
SrMnSb2[103] $297$[104] $189$ $283.5$ $2$ $4.5$
TbFeO3[105] $681$[105] $456$ $638.4$ $5/2$ $6.2$
TmFeO3[106] $635$[107] $442$ $618.8$ $5/2$ $2.6$
YbFeO3[108] $627$[109] $440$ $616$ $5/2$ $1.8$
YbMnBi2[13] $290$[13] $203$ $284.2$ $5/2$ $2.0$
YbMnSb2[12] $345$[12] $266$ $372.4$ $5/2$ $7.9$
YFeO3[110] $644.5$[111] $445$ $623$ $5/2$ $3.3$
YMnO3[112] $70$[112] $40$ $60$ $2$ $14.3$
YVO3[113] $118$[114] $59$ $118$ $1$ $0.0$